OpenStudy (anonymous):

rex mixed x liters of fruit syrup with y liters of water to make a fruit punch. The number of liters of water he mixed was 6 more than twice the number of syrup he mixed. the total number of liters of syrup and water he mixed was 7 liters. Which graph best shows the number of liters of fruit syrup and water that rex made the fruit punch??

OpenStudy (anonymous):

one moment I must get the graphs so you can see

OpenStudy (anonymous):

Start with making the equation. Can you do that? What do you get if you do?

OpenStudy (anonymous):

my first equation was y=6+2x, x+y=7

OpenStudy (anonymous):

Yes, that's what I got as well. Now, do you know what either of those graphs would look like? Start with, in slope-intercept form (Y=mx + b) What is the slope and what is the intercept?

OpenStudy (anonymous):

the m is the slope

OpenStudy (anonymous):

which is 2

OpenStudy (anonymous):

and the other slope is 1?

OpenStudy (anonymous):

Right. So you'll know the slope of the line you're looking for, but do you know how to find the intercept, that's probably the easiest part for you to use to rule out your graphs first.

OpenStudy (anonymous):

the y intercept is 6?

OpenStudy (anonymous):

im not sure which graph will fit